# SLANG: New Concept Comprehension of Large Language Models

### v4

@@ Abstract
[…]
@@ 1 Introduction
Recently, the language evolution has been accelerated by the online community, which has introduced new dimensions to linguistic shifts (Varis and van Nuenen, 2017; Firth et al., 2019; Hammarström, 2016). These rapid changes in language poses serious challenges to the large language models (LLMs) on understanding the newly emerging concepts (Yang et al., 2023; Sun et al., 2021).
Generally, LLMs are trained on static data (Brown et al., 2020), which limits their adaptivity to the dynamic and ever-evolving nature of human language. This limitation is particularly pronounced in the context of digital communication, where new forms of expression and concepts emerge at an unprecedented pace (Sun et al., 2021). Hence, it is essential for LLMs to to understand linguistic shifts and new concepts without constant updates or external data.
@@
FOCUS employs causal inference to enhance models’ comprehension of new concepts within evolving linguistic contexts. By analyzing causal relationships in language, FOCUS advances models’ predictive capabilities beyond traditional correlation-based learning. This method allows for a nuanced grasp of language dynamics, improving models’ adaptability and effectiveness in applications requiring deep understanding of language use. FOCUS significantly enhanced performance in language model comprehension, demonstrating superior precision and adaptability. With GPT-4, FOCUS achieved an $\text{F}_{1}$ score of 0.4532, precision of 0.4598, and recall of 0.4551, alongside an accuracy of 84.9%, outperforming previous methods in comprehension and adaptability.
@@ 2 SLANG
We introduce SLANG in response to rapidly evolving language. SLANG benchmark evaluates LLMs’ capability to interpret the dynamic landscape of user-generated new concepts. It uniquely features factual and counterfactual datasets, each crucial for gauging LLM adaptability. We detail SLANG’s problem definition, dataset construction, and evaluation metrics in the following subsections.
2.1 Problem
Diving deep into the digital deluge, we encounter the fluid and fleeting world of internet parlance, where slang and idiomatic expressions defy static definitions and demand dynamic interpretation by LLMs. Our journey unfolds with the task of capturing the essence of this linguistic fluidity, mathematically expressed as the challenge of estimating $P(Y|W,X;\mathcal{M})$. This formulation seeks to unravel the likelihood of an explanation $Y$ emerging from a phrase $W$ within its context $X$, as interpreted by the model $\mathcal{M}$.
@@ 2.2 Preprocessing
The content filtering of our dataset consists of several steps, with the foremost being the assurance of content novelty. Many UrbanDictionary phrases are now common, potentially pre-included in GPT’s training data. Terms like ghosting or flex illustrate this shift to general usage. To avoid redundancy in LLM predictions, we focused on ensuring that the phrases and definitions in our dataset were new to the LLMs. To solve this, we factored in the official knowledge cut-off date of gpt-4-0613, which is April 2021 as detailed in the OpenAI documentation [^2] https://platform.openai.com/docs/models. To ensure the novelty and relevance of our dataset, we selected concepts that were added to UrbanDictionary after January 2022. This temporal gap was strategically chosen to include recent phrases that may have gained popularity online after the cut-off date but were not yet recorded in UrbanDictionary.
Another essential aspect of content filtering is to analyze user-generated ratings, which we visualize in Figure 2. Comparative histograms in the figure illustrate the distribution of upvotes (left) and absolute upvotes (right) across dataset entries before and after our cleaning process. This stringent data cleaning was crucial in refining the dataset to ensure the quality and reliability of the entries selected for our research.
Additionally, we employed gpt-4-0613 for thorough filtration, ensuring terms were not already memorized by GPT, thus maintaining our dataset’s novelty. We also removed inappropriate content, preserving academic integrity and ensuring quality.
@@
Acknowledging the informal nature of the original dataset, we transformed the data into a uniform, formal dictionary format. This transformation involves simplifying the explanations and examples to make them universally understandable while preserving their original online meaning. The transformation process adhered to a custom-designed template, ensuring consistency and clarity across the concepts. We also enriched each explanation with four synonym-based variants, ensuring the capture of diverse potential responses. When evaluating predictions, we compare the model’s output with all five versions of the transformed explanations (the original plus four synonym-augmented variants) and select the result that most closely matches. This method ensures a more comprehensive and accurate assessment of the model’s capability to interpret and reproduce the factual content.
@@
We aim to assess the adaptability and interpretative depth of LLMs in semantic decoding and contextual understanding by preserving original meanings from the factual dataset while introducing new phrases. This approach maintains the semantic substance but shifts the linguistic form. Our strategy involves creating novel phrases that are linguistically distinct yet semantically congruent with their real-world counterparts. Crafting these phrases requires a creative and analytically rigorous approach, leveraging the rich contextual foundation of the original dataset. These new phrases encapsulate the same ideas as the original meanings but are expressed through a different linguistic perspective. Balancing inventive language use with logical linguistic norms is essential to ensure the new phrases are plausible and coherent. Upon generation, these new phrases are paired with their corresponding original meanings, creating a dataset that contrasts novel linguistic expressions against established semantic structures. This pairing allows for a multifaceted analysis of LLMs’ language processing and comprehension, particularly their ability to separate linguistic form from semantic substance. This methodology sheds light on the nuanced capabilities of LLMs in interpreting language beyond mere recognition, delving into semantics and contextual linguistics.
@@ 2.3 Metrics
For this task, we employed traditional metrics like $\text{F}_{1}$ score, recall, and precision Yang et al. (2018), and added BLEU (3-gram) Papineni et al. (2002) and ROUGE Lin (2004) for stricter quality checks. Considering that language models might output synonymous interpretations with varied wording, we also incorporated sentence similarity measures such as sentence-level similarity and SimCSE Gao et al. (2021), and sentence similarity Team (2023) as Similarity by calculating the cosine similarity of the embeddings after the sentences are input into the model all-mpnet-base-v2 (Team, 2023; Song et al., 2020). A positive sample was considered for accuracy calculation if its SimCSE score exceeded 0.7. We also generated five lexically varied yet syntactically and semantically identical interpretations for each dataset entry. To determine the final metric, we selected the interpretation with the highest BLEU score.
@@ 3.1 Causal Analysis
Structural Causal Models (SCMs) serve as vital tools for elucidating the relationships and influence pathways among variables. We propose a simplified SCM (see Figure 3) to delineate the interpretative processes of an LLM when confronted with novel phrases within their context. IIn this model, users supply both the phrase $W$ and its context $X$, which are then inputted into the LLM, represented by the direct links $X\rightarrow\mathcal{M}\rightarrow Y$ and $W\rightarrow\mathcal{M}\rightarrow Y$, where $\mathcal{M}$ denotes the LLM and $Y$ is the output explanation. These causal links, free from confounders, capture the logical chain from input to interpretation. This SCM sets the stage for our forthcoming discussion where we reintroduce and scrutinize these confounders, thus laying the groundwork for a comprehensive causal analysis.
@@
Initially, we reference a typical SCM framework (Wang et al., 2023; Wang et al., 2022), assuming our SCM as $S=\{X,E,Y\}$, where $X$ denotes the context or input, and $E$ represents the confounder, encompassing phrases and other entities within the context. From the perspective of human understanding of new phrases, the interpretation $Y$ is derived from the context $X$ and its entities $E$. The confounder $E$ can be extracted from $X$, leading to the relationships $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$ in our model. By applying the $do$-operation Verma and Pearl (1990), which denoted as $do(X)$, we follow the guidelines (Wang et al., 2023; Wang et al., 2022) to conduct a rigorous causal inference, ensuring that the main effects of the textual context are captured without losing entity information. This operation aims to isolate the effect of the context $X$ on the confounder $E$. Consequently, the relationship between $X$ and $E$ undergoes a change, becoming $\tilde{X}\leftarrow\tilde{E}$, where $\tilde{E}$ represents the modified entity, and $\tilde{X}$ denotes the context obtained after the $do$ operation which effectively substitutes the actual entity.
@@
In our SCM, denoted as $S=\{X,W,E,Y,R\}$, the output variable $Y$, as an endogenous variable, is influenced by the input context $X$, the input phrase $W$, other entities $E$, and finally, linguistic factors $R$, which include the linguistic structure, style, theme, and cultural background, crucially shape $X$ and constitute the exogenous variables. The model includes direct paths $X\rightarrow Y$ and $W\rightarrow Y$, indicating the immediate influence of context and phrase on the interpretation. The backdoor path $E\leftarrow X\rightarrow W$ and the indirect path $X\rightarrow E\rightarrow W$ demonstrate the mediated effects. The path $R\rightarrow X$ highlights the exogenous influence of linguistic factors on context. This configuration captures the causal mechanisms underlying natural language understanding, emphasizing the direct, indirect, and mediated relationships among the variables. The equation for the causal effect in this context is as follows:
@@ 3.2 Context-based Causal Intervention
This revised approach and formula ensure a more comprehensive and nuanced analysis of the causal dynamics within the SCM, post-intervention. This approach ensures a more robust and bias-free interpretation within the SCM framework. This equation accounts for the altered relationships in the SCM after $do$-operation. The conditional probabilities and summations are now over the new variables $\tilde{E}$ and $\tilde{W}$, while maintaining the original structure’s intent to adjust for confounding effects and capture the influence of modified entities in the context of $X$.
@@
In the Masked Entity Inquiry (MEI) stage, we mask the phrase $W$ within context $X$ to analyze the meaning $Y$ without $W$’s direct influence. This process, $Y_{MEA},\hat{W}{MEA}=\mathcal{M}(X{\text{masked}},E)$, helps the model suggest synonyms or near-synonyms for the masked phrase. MEI focuses on extracting meaning from broader linguistic factors in $X$ and $E$, reducing bias towards $W$’s literal interpretation and enhancing context-based understanding.
@@
GPT-3.5 Direct Ouyang et al. (2022) 0.2219 0.1541 0.4320 0.0477 0.2410 0.0547 0.1531 0.1943 0.6806 47.6
CoT Wei et al. (2022) 0.3922 0.3585 0.4583 0.2273 0.4181 0.1822 0.3135 0.5292 0.7653 76.4
CauView Wang et al. (2023) 0.3500 0.2487 0.4885 0.1676 0.3777 0.1506 0.2676 0.5612 0.7868 72.0
@@
Table 1: Performance results on the factual dataset.
@@ 4.1 Setup
In our experimental setup, we utilized a subset of 408 samples, filtered from our initial dataset of 7220 concepts, focusing on evaluations over GPT-3.5 (gpt-3.5-turbo-0613) and GPT-4 (gpt-4-0613). These models were selected for their advanced capabilities and the relevance of their knowledge bases, which have a cut-off date prior to April 2021. As we discussed in 2.2, this cut-off is crucial to ensure the novelty and relevancy of our dataset, as it aligns with the temporal frame of our selected UrbanDictionary concepts.
@@ 4.2 Results
Results on factual dataset
Our study meticulously evaluated the performance of various language models on the factual dataset, with a particular focus on interpreting new phrases. Table 1 summarizes these findings. We can see that the FOCUS consistently outperforms other approaches across all metrics. Specifically, for the GPT-3.5 model, FOCUS achieves the highest $\text{F}_{1}$ score of 0.4292, precision of 0.4153, and recall of 0.4524, along with the best BLUE, ROUGE, similarity, SimCSE scores, and accuracy at 84.5%. A similar trend is noted with the GPT-4 model. Here, FOCUS again lead the metrics, notably achieving an $\text{F}_{1}$ score of 0.4446, precision of 0.4280, recall of 0.4714, and an impressive accuracy of 88.2%. These results underscore FOCUS’s superior ability to interpret and contextualize slang and idiomatic expressions from UrbanDictionary, as compared to other existing methodologies.
@@
GPT-3.5 Direct Ouyang et al. (2022) 0.1922 0.1252 0.4497 0.0371 0.2120 0.0477 0.1328 0.1936 0.6051 24.3
CoT Wei et al. (2022) 0.3576 0.3138 0.4407 0.2098 0.3865 0.1713 0.2932 0.4390 0.7232 62.3
CauView Wang et al. (2023) 0.3161 0.2489 0.4471 0.1464 0.3439 0.1297 0.2491 0.4031 0.6954 54.6
@@
Table 2: Performance results on the counterfactual dataset.
@@
Table 3: Results of ablation experiments on GPT-4.
Results on counterfactual dataset
Following the analysis of the factual dataset, we extended our evaluation to the counterfactual dataset, which presents hypothetical scenarios altering real-world language usage. The performance results, as detailed in Table 2, reveal insightful trends. Notably, our FOCUS method again outperforms others in metrics with GPT-4, achieving an $\text{F}_{1}$ score of 0.4532, precision of 0.4598, recall of 0.4551, and an accuracy of 84.9%. This reinforces FOCUS’s adeptness at interpreting not only direct language use but also hypothetical, contextually modified scenarios. In essence, our results from both datasets affirm the robustness and versatility of FOCUS in navigating the multifaceted nature of language, effectively bridging the gap between real and hypothetical language use.

### v6

@@ Abstract
Lingrui Mei1,4  Shenghua Liu1,4thanks: Corresponding author. Yiwei Wang2,3  Baolong Bi1,4  Xueqi Cheng1,4 1CAS Key Laboratory of AI Safety, Institute of Computing Technology, CAS 2University of California, Los Angeles  3University of California, Merced 4University of Chinese Academy of Sciences meilingrui22@mails.ucas.ac.cn wangyw.evan@gmail.com {liushenghua,bibaolong23z,cxq}@ict.ac.cn
@@ 1 Introduction
Recently, language evolution has been accelerated by the online community, which has introduced new dimensions to linguistic shifts (Varis and van Nuenen, 2017; Firth et al., 2019; Hammarström, 2016). These rapid changes in language pose serious challenges to the Large Language Models (LLMs) on understanding the newly emerging concepts (Yang et al., 2023; Sun et al., 2021).
Generally, LLMs are trained on static data (Brown et al., 2020), which limits their adaptivity to the dynamic and ever-evolving nature of human language. This limitation is particularly pronounced in the context of digital communication, where new forms of expression and concepts emerge at an unprecedented pace (Sun et al., 2021). Hence, it is essential for LLMs to understand linguistic shifts and new concepts without constant updates or external data.
@@
Focus employs causal inference to enhance models’ comprehension of new concepts within evolving linguistic contexts. By analyzing causal relationships in language, Focus advances models’ predictive capabilities beyond traditional correlation-based learning. This method allows for a nuanced grasp of language dynamics, improving models’ adaptability and effectiveness in applications requiring deep understanding of language use. Focus significantly enhanced performance in language model comprehension, demonstrating superior precision and adaptability. With Claude 3, Focus achieved an $\text{F}_{1}$ score of 0.4596, precision of 0.4452, and recall of 0.4827, alongside an accuracy of 89.7%, outperforming previous methods in comprehension and adaptability.
@@ 2 Slang
We introduce Slang in response to rapidly evolving language. Slang benchmark evaluates LLMs’ capability to interpret the dynamic landscape of user-generated new concepts. It uniquely features factual and counterfactual datasets, each crucial for gauging LLM adaptability. We detail Slang’s dataset construction, and evaluation metrics in the following subsections.
@@ 2.1 Preprocessing
The content filtering of our dataset consisted of several steps to ensure the quality and novelty of the concepts:
  • Temporal Filtering: Many UrbanDictionary phrases are now common and potentially included in LLMs’ training data. To ensure the novelty of our dataset, we leveraged the knowledge cut-off dates of LLMs, strategically selecting phrases that emerged after these dates. For instance, gpt-4-0613 has a knowledge cut-off date of April 2021, as detailed in the OpenAI documentation [^2] https://platform.openai.com/docs/models, and we selected concepts that were added to UrbanDictionary after January 2022. This temporal gap was strategically chosen to include recent phrases that may have gained popularity online after the cut-off date but were not yet recorded in UrbanDictionary.
  • User Rate Filtering: We analyzed user-generated ratings to refine the dataset. Entries with overwhelmingly negative receptions (more than 80% downvotes) were excluded. Comparative histograms in Figure 2 illustrate the distribution of upvotes (left) and absolute upvotes (right) across dataset entries before and after our cleaning process. This stringent data cleaning was crucial in ensuring the quality and reliability of the entries selected for our research. For details on the validation of user-generated votes, see Appendix E.
  • Removal of Inappropriate Content: Inappropriate content, including NSFW material and hate speech, was removed to preserve academic integrity and ensure quality.
  • Novelty Check: To ensure that the concepts in our dataset were unknown to the LLMs, we employed gpt-4-0613 for thorough filtration based on the method described by Yin et al., 2023. Additionally, due to the existence of models with different cut-off dates, we applied the "needle in a haystack" test (Kamradt, 2023, see Appendix D) to confirm the novelty of the knowledge. This meticulous validation process involved strategically embedding the selected phrases into extensive corpora and evaluating the models’ capability to extract them.
After the temporal filtering step, we started with 7220 concepts. The subsequent steps filtered out 5463, 1328, and 21 samples respectively, resulting in a final dataset of 408 usable new concepts.
@@
Following the above filtering steps, we obtained our factual dataset. Acknowledging the informal nature of the original data, we transformed it into a uniform, formal dictionary format (see Appendix B). This process involved simplifying explanations and examples for universal understanding while preserving their original meanings. We adhered to a custom-designed template for consistency and clarity across all concepts. Each explanation was enriched with four synonym-based variants to capture diverse potential responses.
@@
To further evaluate the ability of LLMs in understanding new concepts, we created the counterfactual dataset. This dataset is derived from the factual dataset by preserving the original phrases while modifying the contexts and explanations. Given the factual dataset $\mathcal{D}_{\text{fact}}$, the generation process is structured as follows:
  • Entity Extraction: We extract the entities $\mathbf{e}$ from each explanation $\mathbf{y}$ in $\mathcal{D}_{\text{fact}}$.
  • Counterfactual Replacement: For each entity $e_{i}\in\mathbf{e}$, we generate a counterfactual entity $e_{i}^{\prime}$ that is conceptually divergent from the original entity $e_{i}$ while retaining the original phrase structure.
  • Context Construction: Based on the counterfactual explanations $y_{i}^{\prime}$, we use GPT-4 to generate new contexts $x_{i}^{\prime}$. Consequently, the counterfactual dataset $\mathcal{D}_{\text{cf}}$ is composed of pairs: $\mathcal{D}_{\text{cf}}=\{(p_{1},x_{1}^{\prime},y_{1}^{\prime}),(p_{2},x_{2}^{\prime},y_{2}^{\prime}),\ldots,(p_{n},x_{n}^{\prime},y_{n}^{\prime})\}$.
This approach ensures that while the original phrases are preserved, the contexts and explanations are transformed to convey entirely different meanings. Consequently, each entry in the counterfactual dataset introduces novel concepts that is distinct from the original dataset, providing unique challenges for the LLMs to interpret and understand.
@@ 2.2 Metrics
For this task, we employed traditional metrics like $\text{F}_{1}$ score, recall, and precision Yang et al. (2018), and added BLEU (3-gram) Papineni et al. (2002) and ROUGE Lin (2004) for stricter quality checks. Considering that language models might output synonymous interpretations with varied wording, we also incorporated sentence similarity measures such as sentence-level similarity and SimCSE Gao et al. (2021), and sentence similarity Team (2023) as Similarity by calculating the cosine similarity of the embeddings of all-mpnet-base-v2 (Team, 2023; Song et al., 2020). A positive sample was considered for accuracy calculation if its SimCSE score exceeded 0.7. We also generated five lexically varied yet syntactically and semantically identical interpretations for each dataset entry. To determine the final metric, we selected the interpretation with the highest BLEU score.
@@ 3.1 Causal Analysis
Structural Causal Models (SCMs) serve as vital tools for elucidating the relationships and influence pathways among variables. We propose a simplified SCM (see Figure 3) to delineate the interpretative processes of an LLM when confronted with novel phrases within their context. In this model, users supply both the phrase $W$ and its context $X$, which are then inputted into the LLM, represented by the direct links $X\rightarrow\mathcal{M}\rightarrow Y$ and $W\rightarrow\mathcal{M}\rightarrow Y$, where $\mathcal{M}$ denotes the LLM and $Y$ is the output explanation. These causal links, free from confounders, capture the logical chain from input to interpretation. This SCM sets the stage for our forthcoming discussion where we reintroduce and scrutinize these confounders, thus laying the groundwork for a comprehensive causal analysis.
@@
Initially, referencing a typical SCM framework (Wang et al., 2023; Wang et al., 2022), we assume SCM $S=\{X,E,Y\}$, where $X$ denotes context/input, and $E$ represents confounder, including phrases and other entities within the context. From the perspective of human understanding of new phrases, the interpretation $Y$ is derived from the context $X$ and its entities $E$. The confounder $E$ can be extracted from $X$, leading to the relationships $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$ in our model. By applying the $do$-operation Verma and Pearl (1990), which denoted as $do(X)$, we follow the guidelines (Wang et al., 2023; Wang et al., 2022) to conduct a rigorous causal inference, ensuring that the main effects of the textual context are captured without losing entity information. This operation aims to isolate the effect of the context $X$ on the confounder $E$. Consequently, the relationship between $X$ and $E$ undergoes a change, becoming $\tilde{X}\leftarrow\tilde{E}$, where $\tilde{E}$ represents the modified entity, and $\tilde{X}$ denotes the context obtained after the $do$ operation which effectively substitutes the actual entity.
@@
In our SCM, denoted as $S=\{X,W,E,Y,R\}$, the output variable $Y$, as an endogenous variable, is influenced by the input context $X$, the input phrase $W$, other entities $E$, and finally, linguistic factors $R$, which include the linguistic structure, style, theme, and cultural background, crucially shape $X$ and constitute the exogenous variables. The model includes direct paths $X\rightarrow Y$ and $W\rightarrow Y$, indicating the immediate influence of context and phrase on the interpretation. The backdoor path $E\leftarrow X\rightarrow W$ and the indirect path $X\rightarrow E\rightarrow W$ demonstrate the mediated effects. The path $R\rightarrow X$ highlights the exogenous influence of linguistic factors on context. The equation for the causal effect in this context is as follows:
@@ 3.2 Context-based Causal Intervention
This revised formula ensure a more comprehensive and nuanced analysis of the causal dynamics within the SCM, post-intervention. This approach ensures a more robust and bias-free interpretation within the SCM framework. This equation accounts for the altered relationships in the SCM after $do$-operation. The conditional probabilities and summations are now over the new variables $\tilde{E}$ and $\tilde{W}$, while maintaining the original structure’s intent to adjust for confounding effects and capture the influence of modified entities in the context $X$.
@@
In the Masked Entity Inquiry (MEI) stage, we mask the phrase $W$ within context $X$ to analyze the meaning $Y$ without $W$’s direct influence. This process, $Y_{MEA},\hat{W}_{MEA}=\mathcal{M}(X_{\text{masked}},E)$, helps the model suggest synonyms or near-synonyms for the masked phrase. MEI focuses on extracting meaning from broader linguistic factors in $X$ and $E$, reducing bias towards $W$’s literal interpretation and enhancing context-based understanding.
@@
Mistral-7B Direct Ouyang et al. (2022) 0.1869 0.2655 0.1684 0.0372 0.2078 0.0481 0.1452 0.1507 0.6084 43.1
CoT Wei et al. (2022) 0.3453 0.3227 0.3706 0.1978 0.3657 0.1602 0.2744 0.4713 0.7132 68.5
CauView Wang et al. (2023) 0.3012 0.2787 0.3384 0.1561 0.3272 0.1336 0.2449 0.4862 0.7127 64.3
Focus (Ours) 0.3703 0.3555 0.3878 0.2493 0.3942 0.1866 0.3011 0.5121 0.7469 76.0
@@
Claude 3 Direct Ouyang et al. (2022) 0.2395 0.3538 0.2171 0.0552 0.2645 0.0673 0.1904 0.1947 0.6714 51.4
CoT Wei et al. (2022) 0.4276 0.4082 0.4606 0.2471 0.4492 0.2014 0.3371 0.5628 0.7948 81.2
CauView Wang et al. (2023) 0.3752 0.3550 0.4229 0.2015 0.4035 0.1707 0.3041 0.5778 0.7932 76.6
Focus (Ours) 0.4596 0.4452 0.4827 0.3264 0.4835 0.2373 0.3729 0.6109 0.8354 89.7
Table 1: Performance results on the factual dataset. For results on more models, see Appendix C.1.
@@ 4.1 Setup
Following the preprocessing method described in Section 2.1, we filtered our initial dataset of 7220 concepts, resulting in 408 new concepts. These evaluations were conducted using Claude 3, GPT-4, Mistral-7B, and other popular models. For detailed experimental setup and additional results for other models, please refer to Appendix A and Appendix C.
@@ 4.2 Results
4.3 Experimental Results
Results on the factual dataset
Table 1 summarizes the findings that the Focus method demonstrates exceptional performance across all models. For GPT-4, Focus achieves the highest $\text{F}_{1}$ score of 0.4446, precision of 0.4280, recall of 0.4714, and an accuracy of 88.2%. Similarly, Claude 3 under Focus secures an $\text{F}_{1}$ score of 0.4596 and an accuracy of 89.7%, while Mistral 7B records an $\text{F}_{1}$ score of 0.3703 with an accuracy of 76.0%. These metrics highlight the effectiveness of the Focus method in enhancing the interpretative capabilities of language models.
@@ 4.3 Experimental Results
Mistral-7B Direct Ouyang et al. (2022) 0.1765 0.2402 0.1461 0.0327 0.1983 0.0396 0.1378 0.1520 0.5488 20.0
CoT Wei et al. (2022) 0.3318 0.3018 0.3741 0.1849 0.3593 0.1594 0.2734 0.3960 0.6692 56.2
CauView Wang et al. (2023) 0.2922 0.2567 0.3416 0.1431 0.3197 0.1233 0.2365 0.3515 0.6309 46.3
Focus (Ours) 0.3935 0.3928 0.4005 0.2498 0.4151 0.1891 0.3199 0.5079 0.7343 77.5
@@
Claude 3 Direct Ouyang et al. (2022) 0.2123 0.3143 0.1779 0.0454 0.2435 0.0496 0.1769 0.1733 0.5909 23.1
CoT Wei et al. (2022) 0.3928 0.3605 0.4405 0.2143 0.4138 0.1867 0.3184 0.4499 0.7398 64.7
CauView Wang et al. (2023) 0.3468 0.3101 0.4062 0.1662 0.3791 0.1490 0.2879 0.4012 0.6941 52.4
Focus (Ours) 0.4636 0.4739 0.4618 0.3132 0.4894 0.2361 0.3867 0.5783 0.8216 86.8
Table 2: Performance results on the counterfactual dataset. For results on more models, see Appendix C.2.
@@
Table 3: Results of ablation experiments on GPT-4. For results on more models, see Appendix C.3.
Results on the counterfactual dataset
Following the analysis of the factual dataset, we extended our evaluation to the counterfactual dataset, which presents hypothetical scenarios altering real-world language usage. As detailed in Table 2, the Focus method outshines other techniques, particularly with GPT-4, achieving an $\text{F}_{1}$ score of 0.4532, precision of 0.4598, recall of 0.4551, and an accuracy of 84.9%. Claude 3 also performs well under Focus, securing an $\text{F}_{1}$ score of 0.4636 and an accuracy of 86.8%. Mistral 7B shows a solid performance, with an $\text{F}_{1}$ score of 0.3935 and an accuracy of 77.5%. These results underscore Focus’s robust ability to navigate the challenges posed by modified linguistic contexts.